OpenStudy (anonymous):

how to find the equation of a circle when given the endpoints and the diameter HELP!

OpenStudy (anonymous):

Please post you question and I will be glad to help you.

OpenStudy (anonymous):

your*

OpenStudy (anonymous):

for example determine the equation of a circle : the line segment with endpoints E(3,0) and F(-3,0) is a diameter of the circle

OpenStudy (amistre64):

\[(x-Cx)^2+(y-Cy)^2=r^2\]

OpenStudy (amistre64):

really? what point do you see that is the center of 3,0 and -3,0?

OpenStudy (anonymous):

0,0?

OpenStudy (amistre64):

yep, and how long is half of that line?

OpenStudy (anonymous):

0,0?

OpenStudy (amistre64):

\[(x-0)^2+(y-0)^2=r^2\] \[x^2+y^2=r^2\] so far

OpenStudy (amistre64):

just need to know how far half the diameter is

OpenStudy (anonymous):

which is r

OpenStudy (amistre64):

correct; so if the diamter measures from -3 to 3.... how much would half of it measure out to be?

OpenStudy (anonymous):

0

OpenStudy (amistre64):

0 is the center of the line; that really doesnt MEASURE anything tho. we want a distance covered

OpenStudy (amistre64):

if you start at -3 and hop along to get to 3; how far do you travel in the process?

OpenStudy (anonymous):

6?

OpenStudy (amistre64):



OpenStudy (amistre64):

yes, 6 is the diameter then; and half would be our radius

OpenStudy (anonymous):

so 3 is the r but how would i do that in formula form ?

OpenStudy (amistre64):

by filling in the spot that says "r" id imagine. \[x^2+y^2=r^2\] \[x^2+y^2=3^2\] \[x^2+y^2=9\]

OpenStudy (anonymous):

oh okay Thanks :D

OpenStudy (amistre64):

yep ...